# Corki: Enabling Real-time Embodied AI Robots via Algorithm-Architecture Co-Design

### v1

@@ 3.1. Baseline Embodied AI Algorithms
RoboFlamingo is comprised of two main components: a vision language model (VLM) (alayrac2022flamingo, ) and a LSTM network (hochreiter1997long, ) named policy head. At every time step $t$, the VLM takes visual observations $F_{t}$ and a language instruction $i$ as input, and outputs vision-language tokens $X_{t}$. The robot actions $a_{t}$ are generated through the policy head using given $X_{t}$ (li2023vision, ).
Fig. 3. RoboFlamgino policy head. The vision-language token is from the LLM inference. The outputs are seven-dimensional variables that represent the movements of the robot in the next time step.
We elaborate on the action generation process in Fig. 3. At each time step $t$, the policy head takes the visual-language tokens $X_{t}$ generated by the LLM as input and goes through a LSTM network. The hidden state $h_{t}$ is then mapped to the 7-DoF action through two MLP heads as shown in Equ. 2:
@@
During inference, the policy head maintains a queue of length 12. If the queue is not full, the policy head will predict the action $a_{t}^{\text{pose}}$, $a_{t}^{\text{gripper}}$ and update the hidden state $h_{t}$ for the next step prediction; once the queue reaches its maximum capacity, the earliest tokens that entered the queue will be replaced by latest ones, then, consistent with the training process, the action of the current step $a_{t}$ is given based on the last 12 vision-language tokens $X_{t-11}\thicksim X_{t}$.
@@ 3.2. Basic Corki Algorithm
We predict the continuous trajectory of the nearest future instead of discrete actions. We use a cubic function to fit the motion trajectory of robotic arms. For all the seven variables we need to predict, we output a trajectory for each one of the first six variables ($r_{x}(t),r_{y}(t),r_{z}(t),r_{\alpha}(t),r_{\beta}(t),r_{\gamma}(t)$), the gripper $g$ is still a binary value. Using $r_{x}(t)$ as an example, the model output will be shown as Equ. 4, $t$ is time.
@@
After we change the model output, there are two ways of designing loss. The first one is directly supervising $a,b,c,d$. The second one is to supervise the trajectory with the ground truth. We go for the second one for two reasons. The first reason is that usually, no dataset provides the $a,b,c,d$ ground truth, which means we need to extract it from the trajectory ground truth first, which accumulates errors. Second, these parameters vary significantly and are not conducive to the learning of the neural network.
@@
Because of our design, the robotic control and vision inputs are decomposed, leading to less information captured by the robots. To mimic this process during training, we intentionally insert fewer images. As shown in Fig. 4, vision-language tokens from $t=2$ to $t=4$ are shed by a mask embedding, similar to existing works such as BERT (devlin2019bert, ).
@@ 3.3. Optimizing Corki Algorithm
In the basic Corki algorithm, the length of the trajectory is fixed all the time. Suppose the prediction interval is set to be 16.5 ms. No shorter or longer trajectory can be taken. However, one of the most significant characteristics of robotic applications is that they usually encounter sudden environmental changes.
@@
We thus provide flexibility in the length of the trajectory that is taken. The prediction length will be used as an upper bound of the length of the actual taken trajectory, and early termination is allowed. Again, assuming the prediction interval is 16.5 ms, the actual trajectory can be from 3.3 ms to 16.5 ms, with a stride of $\Delta t$ (which is 3.3 ms assuming the camera sensor works in a 30 Hz frequency). After the robot’s early termination, the model will predict another trajectory for the 16.5 ms.
Early termination gives us some flexibility, but it may not be enough. The reason is that the accuracy is higher when the actual trajectory length is consistent between training and inference. If the actual trajectory length is 6.6 ms in training, the same length should be taken during inference. Suppose the user wants to change the actual trajectory length. In that case, the only way is to train two models, one for 6.6 ms and one for 9.9 ms, and switch during inference, which is unsurprisingly inconvenient for almost all robotic applications.
Fig. 4. Masked policy head. The tokens in the dotted line are not generated through a LLM but instead a mask embedding.
@@
We identify the adaptive trajectory length using a concept called waypoints. For example, for a given trajectory spanning 16.5 ms, a waypoint is defined as a point on the trajectory every 3.3 ms or each time step. In Fig. 5, point $A$ is the starting point, points $B$ to $F$ are the waypoints, and point $F$ is the endpoint. Waypoint identification aims to find a waypoint where the robot’s movement is significant. In our case, significant movements are identified as high curvature or changes in the gripper state.
@@
Fig. 5. Waypoints extraction and identification algorithm. The first waypoint with huge movement will be identified and taken as the endpoint of the trajectory.
@@
3.4. Close-loop Control
Till now, Corki is performing open-loop control. The algorithm will produce a trajectory with various lengths, and until the following inference happens, there is no feedback information for the robots. However, open-loop control can lead to lower performance as it easily accumulates errors.
We modify the open-loop feature. During the execution of each trajectory, we randomly send images back before the endpoint of the trajectory. These images are encoded using a convolutional neural network ViT (dosovitskiy2020image, ). The post-encoding features and tokens generated through the LLM are concatenated and used to predict the subsequent trajectory.
@@ 4. Corki Hardware and System Design
This section introduces the Corki hardware. We accelerate the control process to achieve real-time performance. The input of the control module is the trajectory predicted by Corki algorithm, and the output is torque signals that will be used on the motors in each joint of the robots. We call the control framework task space computed torque control (TS-CTC). We first elaborate the control framework (Sec. 4.1), then analyze the bottleneck and propose Corki accelerator (Sec. 4.2). We further propose an effective approximation strategy to improve the control frequency (Sec. 4.3). We finally describe the system pipeline (Sec. 4.4).
@@ 4.1. Task Space Computed Torque Control
The task space computed torque control (TS-CTC) method is widely used in robotics for precise manipulation tasks due to its ability to handle reference inputs in the task space (murray2017mathematical, ). We show the control framework in Fig. 6.
@@
Fig. 6. Task space computed torque control.
The input of TS-CTC has two parts. The first part is the reference trajectory $x_{d}$, the first order derivative $\dot{x}_{d}$ (velocity) and the second order derivative $\ddot{x}_{d}$ (acceleration) of the reference trajectory. The second part is the joint angles $\theta$ and joint angular velocities $\dot{\theta}$ of the robot arm from the sensors. The output is the joint torque $\tau$. We describe the control process in Equ. 6. To achieve smooth robot control, the frequency of generating torques should be at least 100 Hz (yang2023dadu, ; dantec2021whole, ).
@@ 4.2. Corki Hardware
Fig. 7. The results of the key blocks are reusable, as indicated by the arrows. Later stages consume partial results from early stages.
@@
Fig. 8. Hardware architecture for efficiently solving TS-CTC. Blocks in blue are in the format of a dataflow accelerator, and blocks in yellow are customized circuits.
@@
Leveraging the above analysis, our hardware design has two main goals. First, we aim to customize circuits and data pipelines to maximize intermediate data reuse, achieving high parallelization and performance. Second, we focus on customizing on-chip SRAM design to enable single read and write operations during computation, eliminating extra memory accesses.
Fig. 8 shows the Corki architecture, which consists of two parts. The blue blocks form a dataflow accelerator, where all main operators are connected through three FIFOs and a line buffer (LB). This design enables extreme pipelining; for example, the velocity calculation of the first link can start while the pose calculation of the second link begins. The yellow blocks are customized circuits, with the task space mass matrix unit reusing data from the pose unit and the task space bias force unit reusing data from both the velocity unit and the torque unit. There are occasional stalls in the accelerators due to differing latencies between the dataflow accelerator and the customized circuits. A simple micro-controller manages the control flow of the accelerator.
@@ 4.3. Application-specific Approximate Computing
To quantitatively demonstrate our observation, we perform an experiment. We use a 7-DoF Franka Emika Panda robot arm (gaz2019dynamic, ) and monitor the item-wise changes in the mass matrix while slightly adjusting each joint by an angle. For example, we first record all the items in the mass matrix, then change the first joint by 0.1 radians (approximately 6 degrees), 0.3 radians (approximately 17 degrees), and 0.5 radians (approximately 29 degrees), monitoring the changes in the mass matrix. We repeat the same experiments for all the joints in the robot arm.
@@
However, for the joints in the middle of the robot arm, the situation is different. When joint 2 moves, even a change of 6 degrees results in a maximum absolute change in matrix elements of 0.17 (with a maximum relative change of approximately 15.4%). When the motion increases to 29 degrees, the maximum relative change in elements can be as high as 45.2%. The bottom left figure in Fig. 10 shows that when the middle joints undergo movement, the morphology of the robot arm is significantly changed, necessitating the re-computation of all parameters in the control process.
Fig. 9. The maximum difference in the elements of the mass matrix before and after movements in the joint. The experiments are conducted on a Franka Emika Panda robot arm. The movement consists of rotation with angles of 6 degrees, 17 degrees and 29 degrees on all 7 joints.
Fig. 10. The morphology of the Franka Emika Panda robot arm in different configurations. We change joint 1, joint 2 and joint 5 by 29 degrees and show the difference.
@@
We design a simple yet effective approximate computing method to dynamically update the control parameters, reducing the computational costs in the control process. Specifically, given the input $\theta$, we first compute the probability of each matrix (e.g., Jacobian matrix, mass matrix, etc.) needing an update based on an impact factor derived from the angular movement of each joint. In this process, the joints with a small impact on parameter changes have smaller impact factors, while the joints with a large impact on parameter changes have larger impact factors. The probability computation consumes less than 100 FLOPs, which does not affect the final latency.
If the probability of updating a matrix exceeds a certain threshold, the corresponding computation to generate that matrix is performed. Otherwise, the corresponding elements from the previous control cycle are reused. We observe that over 51% of matrix updates can be avoided without any loss in control accuracy.
@@ 4.4. System Pipeline
There are three key components in the system we propose. First, network inference that happens on the server will predict the trajectory. The parameters of the trajectory will be sent to the controller, which is located on the robot. The controller calculates the high frequency actual control signals to enable the robot to move as the trajectory plans, and the robot will move according to the control signals. During the movement of the robot, at random time steps before the trajectory ends, images will be captured by the camera mounted on the robot. These images will be sent back to the server, while the robot continues to finish the rest of the trajectory. Thus, the communication and robotic control can be executed in a parallel way. When the robot reaches the end of the trajectory, it will capture another image and send it back to the server. A new trajectory will be predicted through the LLM inference using this image and previous images.
@@ 5.1. Experimental Setup
We build Corki on the foundation of RoboFlamingo, but our work is extensible to other action-prediction-based embodied AI robots. We implement the algorithm innovation in PyTorch (paszke2019pytorch, ), where the network output predicts a trajectory instead of a discrete action. This predicted trajectory is then fed back into the simulation environments to test the robot’s task completion capabilities.
We use the Calvin (mees2022calvin, ) dataset and software simulation environments, one of the most widely used embodied AI datasets. Calvin includes 34 different tasks with 22994 demonstrations for training and 1000 sequences for testing. We evaluate our algorithm in two different scenarios: Seen scenarios, where the tasks in the testing set are similar but not identical to those encountered during training, and unseen scenarios, which are more challenging as the tasks are entirely new and have not been encountered during training.
@@
The tasks are categorized into five types: moving an object, turning a switch on and off, pushing and pulling a drawer, rotating an object, and lifting an object. We use two metrics to evaluate the algorithm’s accuracy: success rate and average job length. The success rate is the most straightforward metric for quantifying a single task, calculated as the number of successful sequences divided by the total sequences. Given that the embodied AI algorithms are designed to improve robots’ abilities on long-horizon jobs, we further report the accuracy on finishing a job. Each job contains five consecutive tasks. The average job length measures how many tasks the robot can complete within a job, with a maximum of 5.
@@
  • Mean trajectory error. We compare the geographic distance between the predicted trajectory and the ground truth, using root mean square error (RMSE) as the metric. Generally, a smaller RMSE indicates better trajectory prediction.
@@
We evaluate the inference latency and energy consumption using a Nvidia V100 GPU, measuring power with NVML (vonnvml, ). Control latency and power are measured using an Intel 13th generation i7-13700 CPU. We implement Corki hardware on a Xilinx Zynq-7000 SoC ZC706 FPGA (zc706, ) to assess real hardware performance. Additionally, we establish Wi-Fi communication between a 7-DoF Franka Emika Panda robot arm (gaz2019dynamic, ) and our server to measure communication latency.

### v5

@@ 3.1. Baseline Embodied AI Algorithms
RoboFlamingo is comprised of two main components: a vision language model (VLM) (Alayrac et al., 2022) and an LSTM network (Hochreiter and Schmidhuber, 1997) named policy head. At every time step $t$, the VLM takes visual observations $F_{t}$ and a language instruction $i$ as input and outputs vision-language tokens $X_{t}$. The robot actions $a_{t}$ are generated through the policy head using given $X_{t}$ (Li et al., 2024).
Figure 3. RoboFlamingo policy head. The vision-language token is from the LLM inference. The outputs are seven-dimensional variables representing the robot’s movements in the next time step.
We elaborate on the action generation process in Fig. 3. At each time step $t$, the policy head takes the visual-language tokens $X_{t}$ generated by the LLM as input and goes through an LSTM network. The hidden state $h_{t}$ is then mapped to the 7-DoF action through two MLP heads, as shown in Equ. 2:
@@
During inference, the policy head maintains a queue of length 12. If the queue is not full, the policy head will predict the action $a_{t}^{\text{pose}}$, $a_{t}^{\text{gripper}}$ and update the hidden state $h_{t}$ for the next step prediction; once the queue reaches its maximum capacity, the earliest tokens that entered the queue will be replaced by the latest ones, then, consistent with the training process, the action of the current step $a_{t}$ is given based on the last 12 vision-language tokens $X_{t-11}\thicksim X_{t}$.
@@ 3.2. Basic Corki Algorithm
We predict the continuous trajectory of the nearest future instead of discrete actions. We use a cubic function to fit the motion trajectory of robotic arms. For all the seven variables we need to predict, we output a trajectory for each one of the first six variables ($r_{x}(t),r_{y}(t),r_{z}(t),r_{\alpha}(t),r_{\beta}(t),r_{\gamma}(t)$), the gripper $g$ is still a binary value. Using $r_{x}(t)$ as an example, the model output will be shown as Equ. 4, where $t$ represents time.
@@
We employ the cubic function as the trajectory function for two key reasons. First, in robot trajectory planning, the cubic function inherently captures changes in velocity and acceleration since the derivative and second derivative of the cubic function are continuous, effectively modelling the dynamics of real-world motion. Second, they help mitigate overfitting and enhance robustness to noise.
@@
After we change the model output, there are two ways of designing loss. The first one directly supervises $a,b,c,d$. The second one is to supervise the trajectory with the ground truth. We go for the second one for two reasons. The first reason is that almost no dataset provides the $a,b,c,d$ ground truth, so we need to extract it from the trajectory ground truth first, accumulating errors. Second, these parameters vary significantly and are not conducive to the neural network’s learning process.
@@
Because of our design, the robotic control and vision inputs are decomposed, leading to less information captured by the robots. To mimic this process during training, we intentionally insert fewer images. As shown in Fig. 4, vision-language tokens from $t=2$ to $t=4$ are shed by a mask embedding, similar to existing works such as BERT (Kenton and Toutanova, 2019).
@@ 3.3. Optimizing Corki Algorithm
In the basic Corki algorithm, the length of the trajectory is fixed all the time. Suppose the prediction interval is set to be 165 ms. No shorter or longer trajectory can be taken. However, one of the most significant characteristics of robotic applications is that they usually encounter sudden environmental changes.
@@
We thus provide flexibility in the length of the trajectory that is taken. The prediction length will be used as an upper bound of the length of the actual taken trajectory, and early termination is allowed. Again, assuming the prediction interval is 165 ms, the actual trajectory can be from 33 ms to 165 ms, with a stride of $\Delta t$ (which is 33 ms, assuming the camera sensor works in a 30 Hz frequency). After the robot’s early termination, the model will predict another trajectory for the 165 ms.
Early termination gives us some flexibility, but it may not be enough. The reason is that the accuracy is higher when the actual trajectory length is consistent between training and inference. For example, in training, we predict 5 frames, later in testing, although generalizable, fixing the five-step termination shows better accuracy than other step terminations (like 3 or 4). If the actual trajectory length is 66 ms in training, the same length should be taken during inference. Suppose the user wants to change the actual trajectory length. In that case, the only way is to train two models, one for 66 ms and one for 99 ms, and switch during inference, which is unsurprisingly inconvenient for almost all robotic applications.
Figure 4. Masked policy head. The tokens in the dotted line are not generated through a LLM but instead a mask embedding.
@@
We identify the adaptive trajectory length using a concept called waypoints. For example, for a given trajectory spanning 165 ms, a waypoint is defined as a point on the trajectory every 33 ms or each time step. In Fig. 5, point $A$ is the starting point, points $B$ to $F$ are the waypoints, and point $F$ is the endpoint. Waypoint identification aims to find a waypoint where the robot’s movement is significant. In our case, significant movements are identified as high curvature or changes in the gripper state. The reason is that high curvature suggests a rapid change in the robot’s intended motion, which is often necessary to react to new obstacles, adjusted goals, or other unexpected events. By terminating the current trajectory at a point of high curvature, we allow the system to re-evaluate the situation with updated sensor information and generate a new trajectory, enhancing the robot’s ability to adapt to dynamic environments.
@@
Figure 5. Waypoints extraction and identification algorithm. The first waypoint with huge movement will be identified and taken as the endpoint of the trajectory.
@@
3.4. Close-loop Feature
Till now, Corki lacks the ability to sense environmental changes during the trajectory prediction process. The algorithm generates trajectories, with no feedback until the next inference, However, open-loop control can lead to potential error accumulation in robotic manipulation tasks, compared to close-loop control.
We thus introduce close-loop features. During the execution of each trajectory, we randomly send images back before the endpoint of the trajectory. These images are encoded using an encoder network ViT (Dosovitskiy et al., 2020). The post-encoding close-loop features and tokens generated through the LLM are concatenated and used to predict the subsequent trajectory. These features will supervise the step size, and help Corki adjust the predictions in future iterations, accounting for potential changes in the environment.
@@ 4. Corki Hardware and System Design
This section introduces the Corki hardware. We accelerate the control process to achieve real-time performance. The input of the control module is the trajectory predicted by the Corki algorithm, and the output is torque signals that will be used on the motors in each joint of the robots. The control framework we build our accelerator on is widely used task space computed torque control (TS-CTC) (Murray et al., 2017). While the techniques in this section are broadly applicable beyond LLM settings, our accelerator is specifically designed to rapidly convert trajectories into control signals, which is crucial for real-time performance and central to our contributions. We first elaborate on the control framework (Sec. 4.1), then analyze the bottleneck and propose the Corki accelerator (Sec. 4.2). We further propose an effective approximation strategy to improve the control frequency (Sec. 4.3). We finally describe the system pipeline (Sec. 4.4).
@@ 4.1. Task Space Computed Torque Control
The task space computed torque control (TS-CTC) method is widely used in robotics for precise manipulation tasks due to its ability to handle reference inputs in the task space (Murray et al., 2017). We show the control framework in Fig. 6.
@@
Figure 6. Task space computed torque control.
The input of TS-CTC has two parts. The first part is the reference trajectory $x_{d}$, the first order derivative $\dot{x}_{d}$ (velocity) and the second order derivative $\ddot{x}_{d}$ (acceleration) of the reference trajectory. The second part is the joint angles $\theta$ and joint angular velocities $\dot{\theta}$ of the robot arm from the sensors. The output is the joint torque $\tau$. We describe the control process in Equ. 6. Real-time robotic control, especially for smooth manipulation, necessitates high-frequency updates to joint torques. Generating and applying these torques at a frequency of 100 Hz is crucial for responsiveness and avoids jerky motion (Yang et al., 2023; Dantec et al., 2021).
@@ 4.2. Corki Hardware
Figure 7. The results of the key blocks are reusable, as indicated by the arrows. Later stages consume partial results from early stages.
@@
Figure 8. Hardware architecture for efficiently solving TS-CTC. Blocks in blue are in the format of a dataflow accelerator, and blocks in yellow are customized circuits. The gear indicates that the block retains the use of previously obtained results when in approximate computation mode.
@@
Leveraging the above analysis, our hardware design has two goals. First, we aim to customize circuits and data pipelines to maximize intermediate data reuse, achieving high parallelization and performance. Second, we focus on customizing on-chip SRAM design to enable single read and write operations during computation, eliminating extra memory accesses.
Fig. 8 shows the Corki architecture, which consists of two parts. The blue blocks form a dataflow accelerator, where all main operators are connected through three FIFOs and a line buffer (LB). The yellow blocks are customized circuits. A simple micro-controller manages the control flow of the accelerator.
Data Reuse Strategy.
According to Fig. 7, if each module independently computed its results, there would be significant redundant computation. For instance, the forward kinematics block calculates the pose of each link, which is also needed in the Jacobian and mass matrix computations. To mitigate this, we eliminate redundant calculations by centralizing shared computations, such as link poses, velocities, accelerations, and forces, as illustrated in the data flow accelerator of Fig. 8. Instead of re-computing values for each module, Corki hardware computes these quantities once and shares them across relevant computations. For example, the task space mass matrix unit utilizes precomputed poses and Jacobian matrices, while the task space bias force unit reuses precomputed torque and mass matrices, as shown in the customized circuit of Fig. 8. Since the Jacobian matrix and its transpose are required multiple times during a single control computation cycle, and given that the Jacobian matrix is small (typically at most $6\times 7$ in robotic manipulation tasks (Murray et al., 2017)), we allocate a separate memory copy for the Jacobian transpose. This avoids irregular memory access patterns that could otherwise lead to access conflicts, ensuring more efficient processing. Experimental results demonstrate that using our data reuse strategy results in a $54.0\%$ reduction in total latency.
Pipeline Design.
In the data flow accelerator, the computation of pose, velocity, acceleration, and force must follow a sequential order due to physical constraints, e.g., velocity is the derivative of position and acceleration is the derivative of velocity. Thus, computing the acceleration of a link requires knowing its pose and velocity. However, this dependency exists only within the same link. We find that computations for different links can proceed in parallel. For example, while computing link 1’s force, we can simultaneously compute link 2’s acceleration and link 3’s velocity. This pipelined design significantly reduces the delay for computing forces and torques across all links. Experimental results show that our pipelined design results in a further $69.6\%$ reduction in total latency based on the use of the data reuse strategy. When compared to the baseline hardware implementation without any optimizations (i.e., lacking both the data reuse strategy and pipeline design), the total latency is reduced by $86.0\%$.
Memory Optimization.
@@ 4.3. Application-specific Approximate Computing
To quantitatively demonstrate our observation, we perform an experiment. We use a 7-DoF Franka Emika Panda robot arm (Gaz et al., 2019) and monitor the item-wise changes in the mass matrix while slightly adjusting each joint by an angle. For example, we first record all the items in the mass matrix, then change the first joint by 0.1 radians (approximately 6 degrees), 0.3 radians (approximately 17 degrees), and 0.5 radians (approximately 29 degrees), monitoring the changes in the mass matrix. We repeat the same experiments for all the joints in the robot arm.
@@
However, the situation is different for the joints in the middle of the robot arm. When joint 2 moves, even a change of 6 degrees results in a maximum absolute change in matrix elements of 0.17 (with a maximum relative change of approximately 15.4%). When the motion increases to 29 degrees, the maximum relative change in elements can be as high as 45.2%. The bottom left figure in Fig. 10 shows that when the middle joints undergo movement, the morphology of the robot arm is significantly changed, necessitating the re-computation of all parameters in the control process.
Figure 9. The maximum difference in the elements of the mass matrix before and after movements in the joint. The experiments are conducted on a Franka Emika Panda robot arm. The movement consists of rotation with angles of 6 degrees, 17 degrees and 29 degrees on all 7 joints.
Figure 10. The morphology of the Franka Emika Panda robot arm in different configurations. We change joint 1, joint 2 and joint 5 by 29 degrees and show the difference.
@@
We design a simple yet effective approximate computing method to dynamically update the control parameters, reducing the computational costs in the control process. Whether or not to perform approximate computations is evaluated by the hardware, as shown in Fig. 8. Specifically, given the input $\theta$, we first compute the probability of each matrix (e.g., Jacobian matrix, mass matrix, etc.) needing an update based on an impact factor derived from the angular movement of each joint. In this process, the joints with a small impact on parameter changes have smaller impact factors, while the joints with a large impact on parameter changes have larger impact factors. The probability computation consumes less than 100 FLOPs, which does not affect the final latency.
If the probability of updating a matrix exceeds a certain threshold, the corresponding computation to generate that matrix is performed. Otherwise, the corresponding elements from the previous control cycle are reused. We observe that over 51% of matrix updates can be avoided without any loss in control accuracy, using trajectory error as the metric.
@@ 4.4. System Pipeline
There are three key components in the system we propose. First, network inference that happens on the server will predict the trajectory. The parameters of the trajectory will be sent to the controller, which is located on the robot. The controller calculates the high-frequency actual control signals to enable the robot to move as the trajectory plans, and the robot will move according to the control signals. During the movement of the robot, at random time steps before the trajectory ends, images will be captured by the camera mounted on the robot. These images will be sent back to the server while the robot continues to finish the rest of the trajectory. Thus, communication and robotic control can be executed in parallel. When the robot reaches the end of the trajectory, it will capture another image and send it back to the server. A new trajectory will be predicted through the LLM inference using this image and previous images.
@@ 5.1. Experimental Setup
We build Corki on the foundation of RoboFlamingo (Li et al., 2024), but our work is extensible to other action-prediction-based embodied AI robots. We implement the algorithm innovation in PyTorch (Paszke et al., 2019), where the network output predicts a trajectory instead of a discrete action. This predicted trajectory is then fed back into the simulation environments to test the robot’s task completion capabilities.
We use the Calvin (Mees et al., 2022) dataset and software simulation environments, one of the most widely used embodied AI datasets. Calvin includes 34 different tasks with 22994 demonstrations for training and 1000 sequences for testing. We evaluate our algorithm in two different scenarios: Seen scenarios, where the tasks in the testing set are similar but not identical to those encountered during training, and unseen scenarios, which are more challenging as the tasks are entirely new and have not been encountered during training.
@@
The tasks are categorized into five types: moving an object, turning a switch on and off, pushing and pulling a drawer, rotating an object, and lifting an object. We use two metrics to evaluate the algorithm’s accuracy: success rate and average job length. The success rate is the most straightforward metric for quantifying a single task, calculated as the number of successful sequences divided by the total sequences. Given that the embodied AI algorithms are designed to improve robots’ abilities on long-horizon jobs, we further report the accuracy of finishing a job. Each job contains five consecutive tasks. The average job length measures how many tasks the robot can complete within a job, with a maximum of 5.
@@
  • Mean trajectory error. We compare the geographic distance between the predicted trajectory and the ground truth, using root mean square error (RMSE) as the metric. Generally, a smaller RMSE indicates better robot trajectory.
@@
Inference latency and energy consumption are measured on an Nvidia V100 GPU, with power readings obtained via NVML (von Behren, [n. d.]). We measure latency and energy consumption for control algorithms running on an Intel Core i7-6770HQ CPU. We implement Corki hardware on a Xilinx Zynq-7000 SoC ZC706 FPGA (Xilinx, [n. d.]) to assess real hardware performance. Additionally, we establish Wi-Fi communication between a 7-DoF Franka Emika Panda robot arm (Gaz et al., 2019) and our server to measure communication latency.
The entire evaluation is conducted in this manner. For the baseline, each frame undergoes three stages, LLM inference (real-world latency measured on the server), communication (real-world latency measured through sending actual frames from the robot to the server), control (real-world latency measured on the processor of the robot, which is the Intel Core i7-6770HQ CPU). For Corki, each trajectory undergoes three stages, LLM inference generating trajectory (real-world latency measured on the server), communication (real-world latency measured through sending actual frames from the FPGA to the server), control (real data measured on our FPGA board with real trajectories as inputs). We use synchronized timestamps to accurately measure the communication latency.